### v1

# Hot, Photoionized X-ray Gas in Two Luminous Type 2 Quasars: Chandra-HST Evidence for a Wind-Driven Sequence

Abstract
We present new Chandra/ACIS-S imaging spectroscopy of two luminous type 2 quasars, FIRST J120041.4+314745 (=0.116) and 2MASX J13003807+5454367 (=0.088), and compare their X-ray emission with Hubble Space Telescope [O III]5007 morphologies and kinematics. Both systems show kiloparsec-scale soft X-ray emission. In FIRST J120041, the X-ray morphology is clumpy and closely follows the [O III] structures, with surface-brightness peaks co-spatial with the highest [O III] velocities (600–750 km s-1) and broadest line widths (FWHM1700 km s-1). In 2MASX J130038, the X-rays are more centrally concentrated and only weakly correlated with the largely rotational [O III] kinematics. Spectral modeling indicates that photoionization dominates the soft X-rays in both quasars. The inferred hot-gas reservoirs are substantial, (FIRST J120041) and (2MASX J130038), exceeding the outflowing [O III] masses (volume-normalized) by factors of 4 and 16, respectively. In 2MASX J130038, we also identify a tentative blueshifted Fe XXVI Ly line at keV ( km s-1), consistent with a nascent hot wind confined to the inner few hundred parsecs. Combining these results with a broader sample of twelve type 2 quasars, we argue that luminous quasars evolve along a continuous feedback sequence regulated by the progressive clearing of circumnuclear gas. As AGN radiation and winds pierce through the surrounding medium, systems transition from heavily enshrouded, compact configurations to phases where the X-ray and [O III] components strongly couple and, eventually, to large-scale, energetically dominant outflows. FIRST J120041 and 2MASX J130038 represent two points along this sequence, tracing the emergence and growth of hot winds as primary drivers of quasar-scale feedback.
1 Introduction
Active galactic nuclei (AGNs) have the potential to influence the evolution of galaxies through mass outflows that interact with the surrounding interstellar medium (ISM). These feedback processes may regulate the co-evolution of supermassive black holes (SMBHs) and their host galaxies, helping to establish observed scaling relations such as the - correlation between SMBH mass and bulge velocity dispersion (e.g., gebhardt_relationship_2000; begelman_agn_2004). Despite their potential significance, the physical mechanisms that might drive AGN feedback remain under debate, largely because they are difficult to isolate observationally. The relevant processes operate on different spatial and temporal scales, and many potential mechanisms, such as induction heating, shocks, or outflows (e.g., gaspari_unifying_2017; gaspari_linking_2020; trindade_falcao_hubble_2021; falcone_analysis_2024), leave ambiguous or indirect signatures that make it challenging to identify the dominant driver in any given system.
At low luminosities (bolometric luminosities ), nearby () Seyfert galaxies provide valuable insight through narrow-line region (NLR) outflows. Optical spectroscopy of the [O III] line has been central to mapping NLR kinematics, revealing signatures of AGN-driven winds launched from near the SMBH as well as outflows arising from in situ gas entrained within the host galaxy disk (fischer_determining_2013; crenshaw_feedback_2015; fischer_gemini_2017). Radiative acceleration is often invoked with considerable success to explain these outflows (meena_radiative_2021; meena_investigating_2023; falcone_analysis_2024; tutterow_shape_2025), yet the dynamics and origin of NLR winds are still not fully understood (e.g., gaspari_unifying_2017; gaspari_linking_2020).
At higher luminosities, type 2 quasars (QSO2s) provide a unique window into feedback operating on kiloparsec scales. Using Hubble Space Telescope (HST) [O III] imaging and Space Telescope Imaging Spectrograph (STIS) spectroscopy, fischer_hubble_2018 showed that local, luminous QSO2s (, ) can host NLRs extending over several kiloparsecs (see also e.g., storchi-bergmann_bipolar_2018), with clear outflow signatures in their inner regions (see also e.g., de_oliveira_gauging_2021). However, they also found that not all extended [O III] NLRs are accompanied by equally extended [O III] outflows. This discrepancy raises a key question: if outflows are radiatively accelerated and the AGN is luminous enough to ionize gas across kiloparsec scales, why do some outflows appear to stall near the nucleus?
One possibility is that hot X-ray winds provide the dominant dynamical driver, entraining and accelerating cooler gas phases such as the [O III]-emitting component (trindade_falcao_hubble_2021). Indeed, ultra-fast outflows (UFOs) detected in X-rays have been linked to galaxy-scale molecular winds in ultra-luminous infrared galaxies (ULIRGs; tombesi_wind_2015), suggesting a physical connection between hot winds and cooler outflowing gas. In nearby Seyferts, Chandra imaging reveal extended soft X-ray emission that is spatially coincident with [O III] structures (e.g., young_chandra_2001; wang_deep_2011; maksym_cheers_2019), while blueshifted X-ray emission lines show velocities comparable to those observed in the optical gas (e.g., kraemer_physical_2015; kallman_census_2014). In some systems, the X-ray outflow extends even beyond the optical NLR, implying that the hot phase may dominate the total energy budget (kraemer_mass_2020).
These correlations persist at higher luminosities. The luminous QSO2 Mrk 34, for example, shows a well-defined [O III] bicone and high-velocity outflows () extending to pc (fischer_hubble_2018). Multi-component modeling of its NLR (revalski_quantifying_2018; revalski_erratum_2019) combined with spatially resolved STIS spectroscopy yielded a peak outflow rate at kpc and peak kinetic luminosity at pc (trindade_falcao_hubble_2021-1). The close morphological agreement between [O III] and soft X-ray emission in Mrk 34 reinforces the idea of a physical coupling between these phases.
To investigate this connection further, we obtained new Chandra/ACIS-S observations of two luminous QSO2s from the fischer_hubble_2018 sample: FIRST J120041.4+314745 (, 2.04 kpc arcsec-1; hereafter FIRST J120041) and 2MASX J13003807+5454367 (, 1.59 kpc arcsec-1; hereafter 2MASX J130038). These objects are the two [O III]5007-brightest ( and 42.7, respectively; trindade_falcao_hubble_2021-1) of the extended sources following Mrk 34 and have the largest deprojected NLRs ( kpc and 4.7 kpc, respectively; fischer_hubble_2018). Despite their similar NLR sizes, their [O III] outflows show strikingly different kinematics: FIRST J120041 hosts high-velocity outflows extending to 1 kpc, whereas 2MASX J130038 shows predominantly rotational motion. Photoionization modeling of their STIS spectra underscores this contrast, with FIRST J120041 showing [O III] mass outflow rates and kinetic luminosities that exceed those of 2MASX J130038 by factors of 210 and 100, respectively (trindade_falcao_hubble_2021-1).
These contrasting properties make the two quasars an ideal testbed for isolating the role of hot, X-ray-emitting gas in driving large-scale ionized outflows. In this work, we combine the sub-arcsecond resolution of Chandra with archival HST imaging and spectroscopy to (1) map the spatially resolved ionization structure of the X-ray-emitting gas, and (2) test for morphological and kinematic correspondences between the X-ray and optical phases. By comparing two quasars with similarly extended optical NLRs but drastically different outflow kinematics and energetics, we aim to assess the physical role of hot winds in AGN feedback and to place these systems within a broader evolutionary sequence of quasar activity.
2 Observations and Data Reduction
This study is based on new Chandra/ACIS-S observations of the QSO2s FIRST J120041 and 2MASX J130038, each with a total effective exposure time of 90 ks (Table 1). All datasets are publicly available through the Chandra Data Collection (CDC) “332” (https://doi.org/10.25574/cdc.332 (catalog doi:10.25574/cdc.332)).
| ObsID | Date | Time | PI | Offset |
| (ks) | (dx, dy) | |||
| 2MASX J13003807+5454367 | ||||
| 26787 | 2024-01-22 | 14.21 | Kraemer | -0.360, -0.566 |
| 27160 | 2023-02-09 | 30.20 | Kraemer | -1.427, -1.314 |
| 27161 | 2024-05-16 | 42.02 | Kraemer | +0.586, -0.238 |
| FIRST J120041.4+314745 | ||||
| 26786 | 2024-01-10 | 57.32 | Kraemer | +1.009, -0.189 |
| 27158 | 2024-05-24 | 25.74 | Kraemer | +1.022, -1.382 |
We retrieved the raw event files from the Chandra Data Archive111https://cda.harvard.edu/chaser/ and reprocessed them with CIAO 4.17222http://cxc.harvard.edu/ciao (fruscione_ciao_2006), applying the latest calibration files through the chandra_repro333https://cxc.cfa.harvard.edu/ciao/ahelp/chandra_repro.html script. Subpixel event repositioning was enabled to fully exploit the angular resolution of Chandra and resolve extended structures on sub-arcsecond scales.
To ensure precise registration among individual observations, we refined the astrometry using the fine_astro444https://cxc.cfa.harvard.edu/ciao/ahelp/fine_astro.html routine, which cross-matches detected X-ray sources with an external reference catalog and applies small linear shifts to the event files. The resulting positional offsets (dx, dy) in units of sky pixels are listed in Table 1 and are consistent with the nominal uncertainties of the Chandra aspect solution. The reprocessed and aligned event files were then merged to produce the final datasets used for the imaging analysis described in Section 4.
Spectra were extracted using the specextract555https://cxc.cfa.harvard.edu/ciao/ahelp/specextract.html task. For FIRST J120041, we used a circular aperture of radius centered at R.A.=12:00:41.4456, decl.=+31:47:46.062, while for 2MASX J130038 we adopted a aperture centered at R.A.=13:00:38.2106, decl.=+54:54:36.739. These apertures encompass the full extent of the detected emission. Background spectra were extracted from nearby, source-free regions with radii of .
The individual spectra and corresponding response files were combined using combine_spectra666https://cxc.cfa.harvard.edu/ciao/ahelp/combine_spectra.html, and the resulting merged spectra were binned to a minimum of 10 counts per bin. All spectral fitting was performed in the rest-frame 0.3-8 keV energy range.
3 Spectral Properties of the X-ray Emission
3.1 Phenomenological Models
We first model the spectra using phenomenological, multi-component fits implemented in Sherpa777https://cxc.cfa.harvard.edu/sherpa/. These models are intended to identify the main continuum components and locate prominent emission features without imposing strong physical assumptions (e.g., trindade_falcao_deep_2023).
The baseline model consists of an absorption component (xstbabs) allowed to vary above the Galactic value888calculated using the HEASARC tool: https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl ( cm-2 for FIRST J120041, and cm-2 for 2MASX J130038), combined with a series of redshifted Gaussian lines (xszgauss) modeling discrete emission features. When required by the data, we include a hard X-ray continuum component 3 keV, modeled either as a simple power law (xspowerlaw) or as a physically motivated reflection component using MYTorus999https://www.mytorus.com/ (murphy_x-ray_2009). In the MYTorus configuration, the scattered continuum (myts) and fluorescent line (mytl) components were linked to a common column density, with the inclination fixed at 90∘. Owing to limited photon statistics, the photon index is fixed at for all MYTorus and xspowerlaw components.
Figure 1 shows the best-fit phenomenological models, and the resulting parameters are summarized in Table 2.
| Component | Physical Interpretation | Parameter | Value | Identification |
| FIRST J120041.4+314745 | ||||
| Phenomenological fit (Cstat/d.o.f. = 2.92) | ||||
| xstbabs | Foreground absorption | |||
| xszgauss | Emission Line | O VIII (blended) | ||
| norm | ||||
| xszgauss | Emission Line | Mg XI (blended) | ||
| norm | ||||
| xszgauss | Emission Line | Si XIII (blended) | ||
| norm | ||||
| xspowerlaw | AGN Continuum | norm | ||
| 2MASX J13003807+5454367 | ||||
| Phenomenological fit (Cstat/d.o.f. = 0.76) | ||||
| xstbabs | Foreground absorption | |||
| xszgauss | Emission Line | Mg XI (blended) | ||
| norm | ||||
| xszgauss | Emission Line | Si XIII (blended) | ||
| norm | ||||
| xszgauss | Emission Line | S XV (blended) | ||
| norm | ||||
| xszgauss | Emission Line | S XV | ||
| norm | ||||
| xszgauss | Emission Line | Fe XXVI | ||
| norm | ||||
| myts | Scattered AGN Continuum (torus) | |||
| norm | ||||
| mytl | Fluorescence Lines (torus) | norm | ||
- FIRST J120041
The spectrum of this source is dominated by soft X-ray emission below 2 keV. An absorbed power law gives a good representation of the continuum, while additional Gaussian components are required at keV (O VIII, likely blended with O VII), keV (Mg XI, likely blended with Mg XII/Fe L) and keV (Si XIII, possibly blended with Si XIV). The best-fit line-of-sight absorption is cm-2 (best-fit model Cstat/d.o.f. = 2.92).
- 2MASX J130038
The spectrum shows a pronounced hump between 3-6 keV, along with distinct emission features at both lower and higher energies. The best-fit model (Cstat/d.o.f. = 0.76) requires a myts scattered continuum, its associated mytl fluorescent line component, and an additional Gaussian at keV. Below 3 keV, several Gaussian lines are detected at keV (Mg XI, possibly blended with Mg XII), keV (Si XIII/Si XIV), keV (S XV/S XVI), and keV (S XV). The 7.14 keV feature is discussed further in Section 6.2 as a candidate blueshifted Fe XXVI Ly line. The best-fit line-of-sight absorption is cm-2.
3.2 Physically Motivated Models
To investigate the physical origin of the X-ray emission, we next fit models composed of photoionized and, when necessary, thermal plasma components. The photoionized emission is modeled using Cloudy (chatzikos_2023_2023), while thermal emission is tested with Apec (foster_updated_2012). All fits adopt the phenomenological continuum from Section 3.1, with the intrinsic absorbing column allowed to vary above the Galactic value.
The Cloudy grids assume a broken power-law spectral energy distribution (SED) of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5-10 keV, and from 10-100 keV, adopting 1.4 solar abundances (trindade_falcao_hubble_2021-1). The final grids span in steps of 0.25 dex and in 0.1 dex steps. The spectra are exported as XSPEC-style additive tables (porter_cloudyxspec_2006) and fitted using Sherpa.
A physical requirement of the photoionization models is that the emitting slab must be geometrically thin at the distance of the gas. Specifically, the slab thickness should satisfy (e.g., kraemer_physical_2015). This ensures that the cloud does not exceed either the amount of material available at that radius or the maximum emitting surface implied by the morphology. In practice, this condition introduces a coupled constraint between the ionization parameter and the allowed range of column densities : for a given , only certain are physically plausible.
Because varies during the fitting process, the upper limit on the column density, cannot be imposed as a single fixed boundary. Instead, we implement a dynamic constraint in Sherpa that updates the maximum allowed column density at each iteration of the fit. The constraint has the general form
motivated by the condition and calibrated to the geometry of each region. In Sherpa, this relation is enforced by a helper function that updates the parameter bounds for the table model. By updating the allowed range dynamically, the fit explores only the region of parameter space that corresponds to physically realizable clouds at the measured radial distances. This produces stable and self-consistent fits while preventing artificial degeneracies between and .
The resulting best-fit physical models are presented in Figure 2, and their parameter values in Table 3.
- FIRST J120041
The spectrum is best described (Cstat/d.o.f. = 1.73) by two photoionized components: one medium-ionization () and medium-column ( cm-2), and one high-ionization () and low-column ( cm-2) component. No additional continuum is required to reproduce the data.
- 2MASX J130038
The best-fit model (Cstat/d.o.f. = 0.76) includes a scattered continuum (myts), the Gaussian line at 7.14 keV, and one medium-ionization (), medium-column ( cm-2) photoionized component dominating the soft band.
Adding a thermal plasma component does not significantly improve either fit, indicating that photoionization alone adequately accounts for the observed X-ray spectra in both quasars.
| Component | Parameter | Value |
| FIRST J120041.4+314745 | ||
| 2 photoionized components (Cstat/d.o.f. = 1.73) | ||
| xstbabs | ||
| CLOUDY 1 | ||
| 21.92 | ||
| norm | ||
| CLOUDY 2 | ||
| 19.92 | ||
| norm | ||
| 2MASX J13003807+5454367 | ||
| 1 photoionized component (Cstat/d.o.f. = 0.76) | ||
| xstbabs | ||
| CLOUDY 1 | ||
| norm | ||
4 Morphological Properties of the X-ray Emission
4.1 PSF Modeling
Accurate modeling of the Chandra point spread function (PSF) is essential for distinguishing real extended emission from unresolved nuclear components. For each observation, we simulate the PSF using the Chandra Ray Tracer (ChaRT101010https://cxc.cfa.harvard.edu/ciao/PSFs/chart2/index.html) and projected the rays onto the detector plane with MARX111111https://space.mit.edu/cxc/marx/. We generate ensembles of 1,000 PSF realizations centered at the source position and merge them within the same energy bands adopted for the imaging analysis (Section 4.2).
Following the empirical method of fabbiano_revisiting_2020, each simulated PSF is normalized to the observed counts within an annulus of 0.2′′-0.5′′ radius. This normalization assumes that the observed emission comprises a central point-like component plus any additional extended structure. By testing multiple normalization radii we ensure robust scaling and minimize the risk of over-subtracting flux from the PSF wings.
4.2 Broad- and Narrow-Band Imaging
We construct PSF-subtracted images in both broad and narrow energy bands. The narrow bands are selected to isolate line-dominated regions: 0.8-1.2 keV (Ne IX-X; “Neon band”), 1.2-1.6 keV (Mg XI-XII; “Magnesium band”), and 1.6-2.0 keV (Si XIII-XIV; “Silicon band”). The broad bands are defined as 0.3-3 keV (“soft band”), 3-8 keV (“hard band”), and 0.3-8 keV (“full band”). For 2MASX J130038, where the spectrum shows additional complexity at energies above 3 keV, we also produce images covering the 3-6 keV and Fe K (6-8 keV) bands.
Figures 3 and 4 show the resulting PSF-subtracted images, smoothed with a Gaussian kernel () and displayed with asinh scaling. White contours trace the [O III] emission (fischer_hubble_2018), allowing for direct comparison between the X-ray and optical morphologies.
In Figure 5, we show the X-ray surface brightness radial profiles for both quasars across several narrow energy bands. The profiles were extracted from the un-smoothed, PSF-subtracted Chandra images using 20 concentric annuli centered on each AGN and spanning radii of . To allow for direct comparison among bands, all curves are normalized to the surface brightness of the innermost bin in the full (0.3–8 keV) band.
- FIRST J120041
The full-band (0.3-8 keV) emission extends to (4.1 kpc). The soft X-rays dominate, showing clumpy structures in multiple directions and two distinct peaks are evident: one located 0.2′′ (410 pc) east of the centroid, likely associated with the peaks seen in the Neon and Magnesium band images, and another 0.2′′ (410 pc) southwest of the centroid, consistent with the peak observed in the Silicon band. The soft-band morphology closely follows the [O III] contours, with the brightest X-ray knots coincident with peaks seen in [O III]. The hard-band emission is less extended (, 2.9 kpc) but shows similar correspondence with the inner [O III] structures. In the narrow bands, the Neon band has a low signal-to-noise ratio (15 counts within 3.5′′) but still shows extended emission peaking 0.35′′ (715 pc) northeast of the nucleus. The Magnesium band (28 counts) shows a clumpy feature 0.6′′ (1.2 kpc) southwest of the nucleus, while the Silicon band (25 counts) peaks west of the nucleus, coincident with the brightest [O III] peak.
- 2MASX J130038
The emission extends to (3.2 kpc) but shows a simpler morphology and weaker correspondence with the [O III] contours. The full-band image shows a bright, compact nucleus surrounded by fainter extended emission. The soft band reveals three distinct regions of enhanced emission: one located 0.6′′ northeast of the centroid, another 0.2′′ to the southeast, and the brightest region positioned 0.3′′ southwest of the centroid. The remaining soft X-ray emission appears as faint clumps extending in multiple directions. The hard band is comparably extended () but dominated by a central component. In the narrow bands, the Neon band (12 counts within 2.5′′) is confined within . The Magnesium band (32 counts) shows peaks both above and below the nucleus, with a faint clump 0.9′′ (1.4 kpc) southwest. The Silicon band (28 counts) peaks 0.4′′ (640 pc) southwest, consistent with the soft-band maximum. At higher energies, the 3-6 keV emission resembles the full-band morphology, while the Fe K band (6-6.8 keV) is extended to (1.6 kpc) along the east-west axis. Emission in the 6.8-8 keV band is also extended, oriented northeast-southwest and confined within the inner 0.5′′ (800 pc).
5 X-ray Gas Modeling
5.1 Photoionization Modeling
We construct photoionization grids using Cloudy version C.23 (chatzikos_2023_2023). The predicted line spectrum depends on the shape of the ionizing SED, the gas density (), column density (), distance from the ionizing source (), and the adopted chemical abundances. The models are parameterized by the dimensionless ionization parameter
where is the speed of light and the ionizing photon rate is given by
Following Section 3.2, we adopt an SED of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5–10 keV, and from 10–100 keV, with a high-energy cutoff at 100 keV. The ionizing luminosities are taken from trindade_falcao_hubble_2021-1, based on the NuSTAR analysis of gandhi_nustar_2014: photons s-1 for FIRST J120041 and photons s-1 for 2MASX J130038 (but see Section 5.2). Chemical abundances are set to solar, consistent with our earlier work.
Because the emitting regions lie at distances of tens of parsecs or more, we adopt an open-slab geometry, following previous studies of extended photoionized gas (e.g., kraemer_physical_2015; revalski_quantifying_2018; kraemer_mass_2020; trindade_falcao_hubble_2021-1).
5.2 Modeled Line Luminosities
To convert the observed line fluxes into gas masses, we require estimates of both the emitting volume and the corresponding hydrogen density. For a photoionized cloud at radius , the density is
At a fixed distance from the nucleus, , the only free parameters in Equation 4 are and . For each model in the Cloudy grid, we therefore select a range of values that reproduce the observed emission lines and solve for the physically consistent density. We adopt , the characteristic average radius of the extraction region, for both quasars.
FIRST J120041 shows emission from multiple ionization states, requiring two photoionized components, hereafter LOW and HIGH. For simplicity, we assume both components are co-spatial (i.e., ). In contrast, the spectrum of 2MASX J130038 is well described by a single photoionized component (HIGH), which we treat as the sole contributor to its observed line luminosities.
To identify the best-fitting model at each location, we compare the modeled and observed line ratios for all fractional combinations of LOW and HIGH components; an ideal match yields a ratio of unity. The ionizing luminosities are uncertain by a factor of 4 (trindade_falcao_hubble_2021-1). For FIRST J120041, the published value reproduces the observed luminosities, whereas for 2MASX J130038 a value larger by a factor of 3.8 is required, still within the allowed range.
Physically, the models are constrained by two requirements: (1) the emitting slab thickness, , cannot exceed the available gas column at that radius, and (2) the emitting area cannot exceed the maximum geometric surface subtended by the source (see also Section 3.2).
5.3 Mass of Ionized Gas
We estimate the total mass of the photoionized X-ray–emitting gas by combining the observed line luminosities with Cloudy photoionization models (Section 5.1). Following kraemer_physical_2015; kraemer_mass_2020, the mass of gas at radius is
where is the mean mass per proton, is the proton mass, is the observed line luminosity, and is the modeled line flux. The term in parentheses represents an effective emitting area, ; multiplying this area by yields the total number of particles covering the source, which scales to mass via .
Observed fluxes are measured using srcflux121212https://cxc.cfa.harvard.edu/ciao/ahelp/srcflux.html and converted to luminosities using distances from fischer_hubble_2018. Table 4 lists the observed and modeled line luminosities within for each target.
For FIRST J120041, the masses of the two ionization components are computed separately and summed, . This is achieved by expressing the observed luminosity as
where and are the modeled fluxes of each component. We solve for the emitting areas by fitting the Neon, Magnesium, and Silicon bands simultaneously, using the SciPy differential_evolution131313https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.differential_evolution.html algorithm with geometric constraints as bounds. Uncertainties are estimated by perturbing the observed luminosities within their measurement errors and repeatedly re-optimizing in a Monte Carlo framework.
5.4 Comparison with Optical Ionized [O III] Gas
We now compare the mass of the X-ray–emitting gas to that of the cooler ionized phase traced by [O III]. For simplicity, we assume that each phase is uniformly distributed within its respective extraction region. Under this assumption, a direct comparison requires accounting for the fact that the two phases occupy different volumes. We therefore define a volume-normalized X-ray mass
where and are the emitting volumes of the [O III] and X-ray phases. The ratio thus quantifies the relative contributions of the hot and warm ionized gas under a common geometric framework.
In reality, local gas density is expected to decrease with radius (e.g., revalski_quantifying_2022). Moreover, the [O III] phase, having lower ionization and higher density than the X-ray–emitting gas, requires a correspondingly smaller volume to reproduce its observed luminosity. These effects imply that assuming uniform density and unity filling factor tends to overestimate the emitting volumes, especially for the X-ray gas. Consequently, the volume-normalized mass ratios derived below should be regarded as conservative upper limits, particularly in systems where the X-ray–emitting gas extends over a substantially larger region than the [O III] gas.
- FIRST J120041
Within (2 kpc), the low- and high-ionization X-ray components contain and , respectively. The total X-ray mass is therefore dominated by the low-ionization phase: . The [O III] outflow contains (trindade_falcao_hubble_2021-1) within (1 kpc). After accounting for the different emitting volumes (Equation 7), the resulting volume-normalized mass ratio is .
- 2MASX J130038
Within (1.6 kpc), the X-ray–emitting gas contains . The [O III] outflow contains within (0.16 kpc). After correcting for the different emitting volumes, the resulting volume-normalized mass contrast is .
For comparison, nearby Seyfert galaxies host significantly smaller X-ray gas reservoirs. NGC 1068 contains within the inner pc, combining low- and high-ionization components (kraemer_physical_2015), while NGC 4151 has , within pc (kraemer_mass_2020). In, the X-ray gas masses of FIRST J120041 and 2MASX J130038 exceed these values by factors of 300-800 (2.5-3 dex), though measured within volumes 30-8000 times larger.
| Band | Modeled | ||
| (keV) | Lines | ( erg s-1) | ( erg s-1) |
| FIRST J120041.4+314745 | |||
| 0.8-1.2 | Ne X Ly | ||
| 1.2-1.6 | Mg XI-XII | ||
| 1.6-2.0 | Si XIII-XIV | ||
| 2MASX J13003807+5454367 | |||
| 0.8-1.2 | Ne X Ly | ||
| 1.2-1.6 | Mg XI-XII | ||
| 1.6-2.0 | Si XIII-XIV | ||
| ( cm2) | ( cm2) | () | () | () |
| FIRST J120041.4+314745 | ||||
| 4.5 | 0.31 | |||
| 2MASX J13003807+5454367 | ||||
| - | 4.7 | - | ||
6 The Role of X-ray Gas
6.1 [O III] Gas Kinematics
To investigate whether the X-ray and optical phases are dynamically correlated, we compare the X-ray surface brightness with the kinematics of the ionized [O III] gas (fischer_hubble_2018). In that study, the authors defined three kinematic classes based on the [O III] velocity centroids relative to systemic and line widths, which we adopt here:
AGN-driven outflows: [O III] lines with centroid velocities km s-1 or multiple kinematic components.
Disturbed rotation: Single-component [O III] lines with FWHM km s-1 but centroid velocities km s-1.
Non-disturbed rotation: [O III] lines with centroid velocities km s-1 and FWHM km s-1, consistent with ordered disk rotation (see also bellocchi_vltvimos_2013; ramos_almeida_infrared_2017).
We use the dmextract141414https://cxc.cfa.harvard.edu/ciao/ahelp/dmextract.html task to measure the surface brightness of the 0.3-3 keV soft X-ray emission along rectangular extraction strips aligned with the HST/STIS slit positions from fischer_hubble_2018. Each strip spans in length and is divided into nine bins, matching the angular resolution of Chandra. Figures 6 and 7 show, respectively, the extraction geometry overlaid on the 0.3-3 keV Chandra images (top left) and the STIS slit over the HST [O III] images (top right). The bottom panels compare the spatially resolved X-ray surface brightness profiles (orange) with the [O III] velocity centroids (purple, bottom left) and [O III] FWHM values (green, bottom right) for both quasars.
- FIRST J120041
The soft X-ray surface brightness profile confirms that the emission peaks off-nucleus, with a clear enhancement to the east of the nucleus ( pc). Out to at least kpc in that direction, the X-ray surface brightness remains higher than at the nuclear position. Comparing this distribution to the [O III] outflowing kinematics (purple points above the reference line in the bottom-left panel), we find that regions of enhanced X-ray surface brightness (40 counts arcsec-2) coincide with locations where the [O III] gas shows the highest velocities (-750 km s-1) and the broadest line widths (FWHM up to 1700 km s-1). This spatial correspondence shows that the brightest, hot X-ray-emitting gas is co-spatial with the outflowing [O III] component, suggesting a direct interaction between the two phases. While the [O III] flux is higher in the right side (western) emission knot, the outflow velocities and FWHMs are greater on the left (eastern) side, exactly where the soft X-ray peak happens. This further supports the interpretation that the enhanced X-ray emission pinpoints regions of strong coupling between the hot ionized and the cooler [O III] outflow detected by fischer_hubble_2018.
- 2MASX J130038
The soft X-ray surface brightness profile also peaks off-nucleus, showing enhancements on both sides of the nucleus. Out to 1 kpc in either direction along the slit, the extended emission has higher surface brightness than the nuclear region. Although the per-bin X-ray surface brightness (orange points in Figure 7) is higher than in FIRST J120041, the [O III] kinematics in this quasar (purple points, bottom-left panel) are consistent with ordered rotation, i.e., not-outflowing (fischer_hubble_2018). The centroid velocities remain below the threshold for outflowing gas (purple line), and the FWHM values are generally below the disturbed gas threshold (green line), except for a localized broadened component pc northwest of the nucleus. This broadened region coincides with the base of the linear [O III] structure observed near the nucleus. In the X-ray image, an enhancement appears at the end of this structure (see also Figure 4) and another to the southeast, at the location of the prominent [O III] arc. The combination of high FWHM (250 km s-1) but low centroid velocity (300 km s-1) at these position falls into the “disturbed gas” category of fischer_hubble_2018, suggesting that small-scale turbulence or nascent outflow activity could be arising from the linear [O III] feature. Despite its relatively high X-ray surface brightness, no [O III] outflows are detected in this system, and Figure 7 shows no clear spatial correlation between the X-ray enhancements and the [O III] velocity or line-width distributions.
6.2 A Relativistic Hot Wind?
The spectrum of 2MASX J130038 shows a strong neutral Fe K line at 6.4 keV (EW = 228 eV) together with an emission feature at keV (Figures 1, 2). While the mytl component (shown in red) successfully reproduces both the neutral Fe K line at keV and Fe K at keV, an additional Gaussian component centered at keV is required to adequately model the data (significant at 1.4, or 93% confidence).
If interpreted as blueshifted Fe XXVI Ly line (rest energy 6.97 keV), this feature would trace highly-ionized gas either in the immediate vicinity of the nucleus or within a fast nuclear outflow. The observed energy shift corresponds to a velocity of approximately 7,000-8,000 km s-1, consistent with photoionized winds or disk-driven outflows frequently detected in luminous AGNs (e.g., tombesi_evidence_2010; tombesi_wind_2015).
An alternative identification such as Fe K fluorescence from neutral iron (rest energy 7.06 keV) would require a much larger flux, well above the expected 10-15% of Fe K flux (molendi_iron_2003; bianchi_soft_2006). Moreover, the mytl component already reproduces the expected Fe K line, whose flux is 16% of the fitted Fe K line intensity, as shown in Figures 1 and 2. Emission from Ni K (rest energy 7.47 keV) lies too far from the measured centroid to provide a satisfactory explanation, though this possibility cannot be completely ruled out given the current data.
We therefore interpret the 7.14 keV feature as tentative evidence for Fe XXVI Ly emission arising from a hot, blueshifted nuclear wind. In contrast to the absorption-dominated UFOs commonly observed in quasars (e.g., tombesi_discovery_2010), this candidate outflow is detected in emission, implying a geometry and covering factor that allow significant reprocessing of the central radiation field. Deeper hard X-ray observations will be required to confirm its presence and to better constrain its physical conditions and geometry. We discuss the implications of this putative hot wind in the context of a proposed evolutionary sequence for quasars in Section 7.
7 An Evolutionary Path for Quasars
The QSO2s analyzed by fischer_hubble_2018 span a wide range of kinematic and morphological properties. Their sample includes twelve of the fifteen most luminous type 2 quasars in the reyes_space_2008 catalog with . All targets have erg s-1, placing them in the brightest of the QSO2 population from reyes_space_2008. The corresponding black hole masses lie in the range – (e.g., fischer_hubble_2018). Most sources are radio-quiet, with a minority classified as intermediately radio-loud, indicating that their observed outflows are unlikely to be jet-driven.
All twelve objects host [O III]-emitting NLRs that extend beyond 1 kpc, yet only two show outflows that propagate on similarly large scales. fischer_hubble_2018 argued that the distribution of circumnuclear gas near the nucleus may regulate how far ionizing radiation and winds can penetrate into the NLR: in morphologically compact systems, dense inner material can absorb a significant fraction of the AGN ionizing flux, confining outflows to smaller radii. Expanding on this picture, trindade_falcao_hubble_2021 proposed that X-ray winds may progressively entrain this circumnuclear gas, eventually enabling a transition into a large-scale blow-out phase.
In this work, we propose that these quasars evolve along a continuous sequence regulated by the progressive clearing of circumnuclear material by the AGN. In the earliest, most obscured phase, both ionizing radiation and winds are confined to the nuclear region. As the surrounding medium becomes more porous, low-density channels begin to form, allowing ionizing radiation to reach larger radii and enabling high-ionization winds to emerge from the inner few hundred parsecs. Once the winds fully break through, they can drive large-scale outflows capable of reshaping the host ISM.
The differing propagation timescales of ionizing radiation and mechanical feedback naturally account for systems with extended ionized [O III] emission but no large-scale winds. A wind traveling at km s-1 requires yr to cross 1 kpc, whereas the light-crossing time to the same distance is only yr. Consequently, photoionization can illuminate kiloparsec-scale [O III] structures long before mechanical winds have propagated to comparable radii. This evolutionary framework, summarized schematically in Figure 8, links the diversity of morphologies and kinematics observed by fischer_hubble_2018 to different stages along a continuum sequence of circumnuclear clearing and feedback coupling.
Our Chandra observations support this proposed scenario. The X-ray and [O III] properties of 2MASX J130038 and FIRST J120041 suggest that they occupy different locations along the early portion of this evolutionary sequence. In 2MASX J130038, the X-ray emission is extended but only weakly correlated with the [O III] morphology, while the optical kinematics remain dominated by ordered rotation with localized turbulence near the nucleus. Combined with the tentative detection of a blueshifted Fe XXVI line at km s-1 (Section 6.2), these features point to a system in the onset of a blow-out stage (Stage II; see Figure 8), in which a nascent hot wind is beginning to leak out of the nucleus and interact with the inner circumnuclear medium.
In contrast, FIRST J120041 shows clumpy, extended X-ray emission that spatially correlates with the [O III] bicone. Peaks in X-ray surface brightness coincide with regions of the highest [O III] velocities and line widths, pointing to a morphological and kinematic correspondence between the hot and cooler ionized phases. This correspondence suggests that the X-ray wind has begun to couple dynamically to the [O III]-emitting gas, consistent with a relatively more developed stage of the sequence (Stage III; see Figure 8).
Taken together, these quasars may trace the transition from a kinematically compact, blow-out configuration (Stage II) to the onset of a more extended, early post–blow-out state (Stage III). For clarity, in Figure 8 we illustrate this sequence through four representative “stages,” while Table 6 summarizes how the twelve QSO2s studied by fischer_hubble_2018 may populate these stages. These are labeled as follows:
- Stage I: Pre-Blow-Out
This initial phase represents heavily enshrouded systems in which both the X-ray and [O III] emission are morphologically compact and centrally concentrated. Ionizing radiation and winds remain trapped by dense circumnuclear gas, preventing large-scale outflows to eventually form. In the fischer_hubble_2018 sample, examples may include 2MASX J11001238, 2MASX J14054117, and 2MASX J17135038, which all show compact [O III] morphologies ( kpc; fischer_hubble_2018) but broad nuclear line widths (FWHM1000 km s-1; fischer_hubble_2018) indicative of turbulent but spatially confined gas in the inner region.
- Stage II: Blow-Out
Systems where the AGN ionizing radiation begins to escape through low-density channels, ionizing more extended structures but still lacking large-scale outflows. 2MASX J130038 exemplifies this behavior: its [O III] emission extends to kpc, yet the gas is largely rotational, with only localized disturbances within 160 pc of the nucleus (fischer_hubble_2018). The X-rays are morphologically centrally peaked and weakly correlated with [O III] emission (Figure 4), implying that the hot and cooler [O III] phases remain decoupled. The tentative identification of Fe XXVI emission at 7.14 keV ( km s-1) could point to the onset of a nascent hot wind beginning to interact with the inner circumnuclear gas. Similar Stage II behavior is seen in other systems such as 2MASX J07594101, 2MASX J080252, SDSS J115245, B2 1435, Mrk 477, and 2MASX J16531506, where [O III] ionized emission is extended (3 kpc; fischer_hubble_2018) but the kinematics remain dominated by broad line widths and low velocity centroids (FWHM 600 km s-1, 600 km s-1; fischer_hubble_2018). The large X-ray gas reservoir found in 2MASX J130038 likely reflects the accumulation of material during the first evolutionary phases, when dense circumnuclear gas remained confined close to the AGN. As this gas lies nearer to the ionizing source, it reaches higher ionization states, naturally producing a more massive X-ray–emitting component at the onset of blow-out.
- Stage III: Early Post-Blow-Out
As the wind begins to break through the circumnuclear material, the X-ray outflow expands, producing outflowing [O III] kinematics and the spatially correlated X-ray and optical emission we observe. FIRST J120041 shows bright [O III] emission extending to kpc, with multi-component, high-velocity profiles (-750 km s-1) and broad line widths (FWHM 1700 km s-1; fischer_hubble_2018). The soft X-ray emission is clumpy and structurally complex, showing a strong morphological correspondence with the [O III] gas (Figure 3). Peaks in the soft X-ray brightness are co-spatial with regions of outflowing [O III] gas (Figure 6), suggesting that the X-ray wind is now entraining and driving the optical phase. These properties are consistent with a system in which a blow-out has recently occurred, marking the transition from a confined to an expanding feedback regime.
- Stage IV: Late Post-Blow-Out
The nuclear X-ray winds have now cleared most of the circumnuclear material, driving powerful, large-scale outflows. Mrk 34 exemplifies this phase: its [O III] emission extends in a bicone to kpc, showing broad, high-velocity components ( km s-1, FWHM 1700 km s-1; fischer_hubble_2018). It hosts the highest [O III] mass outflow rate and kinetic luminosity in the sample, suggesting that the dense circumnuclear medium has been largely expelled, allowing the AGN to efficiently couple its energy to the host ISM and sustain galaxy-scale feedback (trindade_falcao_hubble_2021-1).
Taken together, these stages outline a coherent evolutionary pathway for quasars. The sequence from pre-blow-out to late post-blow-out phases reflects the gradual coupling of the X-ray and [O III] ionized gas phases as the AGN clears its environment, transitioning from compact, enshrouded systems to those exhibiting large-scale, energetically dominant outflows. A subsequent stage in this sequence may be represented by (unobscured) type 1 quasars (QSO1s), in which the circumnuclear gas has been completely cleared, giving the AGN a “naked” appearance, an unobscured view of the central engine once feedback has effectively evacuated the surrounding medium.
Supporting this scenario, trindade_falcao_hubble_2024 found that the sizes of [O III]-emitting NLRs in type 1 AGNs (spanning erg s-1) scale with luminosity as , a significantly steeper relation than that found for type 2 AGNs over the same luminosity range, . The combined type 1 + type 2 population follows an intermediate slope of . For the subset of seven QSO1s, trindade_falcao_hubble_2024 additionally found significantly extended radial mass profiles, with the bulk of the [O III]–emitting gas located at radii kpc. Relative to the twelve QSO2s in fischer_hubble_2018, these QSO1s have NLRs larger by roughly a factor of three yet contain smaller total [O III] gas masses, consistent with systems in which feedback has already cleared much of the circumnuclear medium and redistributed the remaining ionized gas to larger scales.
Future analysis of [O III] kinematics in these QSO1s (Trindade Falcão et al., in prep.) will further constrain how feedback operates in this unobscured phase.
| Object | [O III] | FWHMnuclear | Stage | ||||
| Morphology | (kpc) | (kpc) | (km s-1) | (kpc) | (km s-1) | ||
| 2MASX J07594101 | Extended, conical | 3.2 | 0.67 | 315 | 0.67 | 1680 | II – Blow-Out |
| 2MASX J08025293 | Extended, biconical | 3.6 | 0.57 | 700 | 0.89 | 870 | II – Blow-Out |
| Mrk 34 | Extended, biconical | 2.2 | 1.89 | 1100 | 1.89 | 580 | IV – Late Post-Blow-Out |
| 2MASX J11001238 | Compact | 3.0 | 0.68 | 350 | 1.51 | 1780 | I – Pre-Blow-Out |
| SDSS J115245 | Extended, biconical | 3.6 | 0.15 | 300 | 1.23 | 360 | II – Blow-Out |
| FIRST J120041 | Extended, biconical | 6.1 | 1.07 | 450 | 1.59 | 720 | III – Early Post-Blow-Out |
| 2MASX J130038 | Extended, biconical | 4.7 | 0.16 | 100 | 0.16 | 580 | II – Blow-Out |
| 2MASX J14054117 | Compact | 1.0 | 0.33 | 100 | 0.94 | 760 | I – Pre-Blow-Out |
| B2 1435 | Extended, biconical | 3.8 | 0.20 | 200 | 1.74 | 630 | II – Blow-Out |
| Mrk 477 | Extended, conical | 2.7 | 0.54 | 500 | 0.90 | 2040 | II – Blow-Out |
| 2MASX J16531506 | Extended, biconical | 6.50 | 0.57 | 110 | 1.23 | 510 | II – Blow-Out |
| 2MASX J17135038 | Compact | 1.8 | 0.65 | 160 | 0.92 | 1660 | I – Pre-Blow-Out |
8 Summary and Conclusions
We present new Chandra/ACIS-S imaging spectroscopy of two luminous QSO2s from the fischer_hubble_2018 sample, FIRST J120041.4+314745 and 2MASX J13003807+5454367, and compare their X-ray properties with HST [O III] morphologies and kinematics. Our analysis combines PSF modeling and subtraction (Section 4.1), broad- and narrow-band imaging (Section 4.2), phenomenological spectral modeling (Section 3.1), and physically motivated photoionization modeling using Cloudy (Section 3.2, and Section 5). The main results are as follows:
- 1
Both quasars show extended X-ray emission, but with different levels of correlation to the optical phase. In FIRST J120041, the X-rays form clumpy, kiloparsec-scale structures that closely trace the [O III] bicone (Figure 3); peaks in soft X-ray surface brightness coincide with the regions of highest [O III] velocities (600-750 km s-1) and broadest line widths (FWHM up to 1700 km s-1; Figure 6). In contrast, 2MASX J130038 shows centrally concentrated X-ray emission with weak morphological correspondence to the [O III] morphology (Figure 4) and no clear spatial correlation with the largely rotational [O III] kinematics (Figure 7).
- 2
Photoionization dominates the soft X-ray emission in both quasars. For FIRST J120041, two photoionized components ( = 0.5 and 2.5; Table 3) reproduce the observed spectrum without requiring a thermal plasma. For 2MASX J130038, a single medium ionization component ( = 1) combined with a hard reprocessed continuum modeled with MYTorus provides a good fit (Figure 2). Adding a thermal component does not improve the fit quality in either case, implying that the soft X-ray emission is dominated by photoionized gas.
- 3
The inferred hot-gas reservoirs are massive. Within apertures (2 kpc for FIRST J120041 and 1.6 kpc for 2MASX J130038), we estimate and , respectively, based on photoionization modeling (Table 5). Relative to the outflowing [O III] gas (trindade_falcao_hubble_2021-1), the X-ray phase is roughly 4 times more massive (volume-normalized) in FIRST J120041 and 16 times more massive in 2MASX J130038. These results indicate that the hot, photoionized phase can dominate both the mass and energy budgets of the ionized ISM on kiloparsec scales.
- 4
2MASX J130038 shows evidence for a nuclear hot wind. An emission feature at keV is consistent with blueshifted Fe XXVI Ly ( km s-1), suggesting a highly-ionized component near the nucleus. This emission may trace a nascent hot wind beginning to emerge within the inner few hundred parsecs (Section 6.2).
- 5
A continuous evolutionary sequence may connect the two quasars and the broader QSO2 population. Integrating the X-ray and optical morphologies, kinematics, and gas masses, we propose that luminous type 2 quasars evolve along a continuum sequence regulated by the gradual clearing of circumnuclear material by AGN radiation and winds (Section 7; Figure 8). In this framework, 2MASX J130038 corresponds to an earlier phase where a weakly coupled, emerging X-ray wind begins to interact with the ISM (analogous to a “blow-out” configuration), while FIRST J120041 represents a relatively more developed stage where the hot wind actively entrains and drives the [O III] gas (an “early post-blow-out” configuration). More compact, enshrouded systems such as 2MASX J11001238 or 2MASX J17135038 may represent the most buried, confined phase of this sequence, “pre-blow-out”, whereas sources like Mrk 34 illustrate the opposite extreme, a “late post-blow-out” characterized by powerful, large-scale outflows and efficient AGN-ISM coupling (Table 6).
Taken together, these results support a scenario in which hot, photoionized X-ray winds may act as the primary drivers of large-scale ionized outflows in luminous type 2 quasars. Where soft X-rays and [O III] emission are morphologically and kinematically correlated (as in FIRST J120041), the winds may be dynamically coupled to the host ISM; where the correlation is weak (as in 2MASX J130038), the wind remains confined and has yet to fully disrupt the circumnuclear gas.
The large inferred hot-gas masses, orders of magnitude above those found in nearby Seyferts such as NGC 1068 and NGC 4151, suggest that the X-ray phase may carry most of the energy and momentum available for feedback on kiloparsec scales.
Finally, the broader distribution of the fischer_hubble_2018 QSO2 sample across this sequence of clearing and coupling implies that luminous obscured quasars may represent different points along a single evolutionary sequence driven by AGN feedback, likely culminating in unobscured type 1 quasars (trindade_falcao_hubble_2024), once the circumnuclear medium has been fully dispersed and the central engine is directly exposed.